# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 3.2 Governing equations
$\eta_{i}$ phase-field order parameter of grain $i$, $i=1,2,3,...,n$
$C_{\text{V}_{\text{O}}^{..}}$ Normalized local concentration of oxygen vacancy ($C_{\text{V}_{\text{O}}^{..}}=a^{3}c_{\text{V}_{\text{O}}^{..}}$, with $a$ being the lattice constant) -
$C_{\text{dop}}$ Normalized local concentration of dopant, $C_{\text{dop}}=a^{3}c_{\text{dop}}$
$C_{\text{V}_{\text{O}}^{..},\text{b}}$ Normalized concentration of oxygen vacancy in bulk phase, $C_{\text{V}_{\text{O}}^{..},\text{b}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{b}}$
$C_{\text{V}_{\text{O}}^{..},\text{c}}$ Normalized concentration of oxygen vacancy in core region, $C_{\text{V}_{\text{O}}^{..},\text{c}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{c}}$
$C_{\text{dop},\text{b}}$ Normalized concentration of dopant in bulk phase, $C_{\text{dop},\text{b}}=a^{3}c_{\text{dop},\text{b}}$
$C_{\text{dop},\text{c}}$ Normalized concentration of dopant in core region, $C_{\text{dop},\text{c}}=a^{3}c_{\text{dop},\text{c}}$
@@
$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$ Formation energy of oxygen vacancy in the bulk phase. $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}N_{\text{A}}$, with $N_{\text{A}}$ being Avogadro constant $\frac{\text{J}}{\text{mol}}$
$g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ Formation energy of oxygen vacancy in the core region. $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{b}}$ Formation energy of dopant in the bulk phase. $g^{0}_{\text{dop},\text{b}}=\mu_{\text{dop},\text{b}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{c}}$ Formation energy of dopant in the core region. $g^{0}_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
@@
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$ Dimensionless available sites of oxygen vacancy in core region. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}=N_{\text{V}_{\text{O}}^{..},\text{c}}a^{3}$, with $a$ being the lattice constant -
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$ Dimensionless available sites of oxygen vacancy in bulk phase. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}=N_{\text{V}_{\text{O}}^{..},\text{b}}a^{3}$
$\tilde{N}_{\text{dop},\text{c}}$ Dimensionless available sites of dopant in core region. $\tilde{N}_{\text{dop},\text{c}}=N_{\text{dop},\text{c}}a^{3}$
$\tilde{N}_{\text{dop},\text{b}}$ Dimensionless available sites of dopant in bulk phase. $\tilde{N}_{\text{dop},\text{b}}=N_{\text{dop},\text{b}}a^{3}$
@@ 4 Finite element implementation of the phase-field model
In this section, we present the detailed finite element implementation for the defect-chemistry informed phase-field model. According to the model presented in Section 3 and Table 2, we treat the OPs $\eta_{i}$ and the field variables $C_{\text{V}_{\text{O}}^{..}}$, $C_{\text{dop}}$, $C_{\text{V}_{\text{O}}^{..},\text{b}}$, $C_{\text{V}_{\text{O}}^{..},\text{c}}$, $C_{\text{dop},\text{b}}$, $C_{\text{dop},\text{c}}$, $\phi$ as degrees-of-freedom (DOFs). Moreover, to solve Eqs. (35) and (36), we use the mixed finite element formulation and treat the electro-chemical potentials $\mu_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}$ as additional DOFs [44].
The strong forms of the corresponding governing equations have already been given in Eqs. (24), (25) and (33) to (39). According to the FEM [45], we formulate the residuals from the weak forms of these governing equations by introducing corresponding test functions $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}$, $\psi_{C_{\text{dop},\text{c}}}$, $\psi_{\eta_{i}}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}$, $\psi_{C_{\text{dop}}}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}$, $\psi_{\mu_{\text{dop}}}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}$, $\psi_{C_{\text{dop},\text{b}}}$ and $\psi_{\phi}$. Firstly, the residuals for the weak forms of Eqs. (24), (25) and (33) are
@@
Following the Galerkin approach [46], the test functions are discretized as $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}=N_{C_{\text{dop},\text{c}}}^{I}\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}=N_{\eta_{i}}^{I}\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}=N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}=N_{C_{\text{dop}}}\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}=N_{\mu_{\text{V}_{\text{O}}^{..}}}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}=N_{\mu_{\text{dop}}}\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}=N_{C_{\text{dop},\text{b}}}\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}=N_{\phi}\psi_{\phi}^{I}$, where $I$ denotes the node index and Einstein summation convention is used. $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}^{I}$ are the nodal weights, while $N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $N_{C_{\text{dop},\text{c}}}^{I}$, $N_{\eta_{i}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{C_{\text{dop}}}^{I}$, $N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{\mu_{\text{dop}}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $N_{C_{\text{dop},\text{b}}}^{I}$ and $N_{\phi}^{I}$ are the shape functions for the corresponding variables. Then we can write down the discretized residuals as
@@
$R_{\eta_{i}}^{I}&=\int_{V}\frac{N_{\eta_{i}}^{I}}{\Delta t}\text{d}V+L\int_{V}N_{\eta_{i}}^{I}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}N_{\eta_{i},j}^{I}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}N_{\eta_{i}}^{I}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}N_{\eta_{i}}^{I}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V\end{aligned}$ (52)
@@
The linearization of the residuals can be written in the element-level linear formulation, in which the terms in the tangent matrix are calculated by $K_{\xi,\zeta}^{IJ}=\partial R_{\xi}^{I}/\partial\zeta^{J}$, where ${\xi,\zeta}$ indicates individual DOF. The non-zero terms in the tangent matrix are presented in supplementary. The numerical implementation of the proposed phase-field model is carried out in Multiphysics Object-Oriented Simulation Environment (MOOSE) framework [47, 48].
@@ 5 Phase-field simulation results of Fe-doped SrTiO3
We present first in Section 5.1 one-dimensional phase-field simulation results to demonstrate its capability of reproducing the formation of a SCL at equilibrium state. Subsequently, we investigate the influence of solute drag effects induced by the segregation of point defects in the core on the formation of an asymmetric SCL under quasi-equilibrium state in Section 5.2. Finally, we conduct two-dimensional (2D) phase-field simulations to study the microstructure evolution process in STO polycrystalline material during sintering and analyze the resulting grain size distribution in Section 5.3.
Note that in the paper we ignore the GB orientation dependency or other anisotropy of the segregation energies $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$, for we intend to reveal the most fundamental mechanisms of SCL formation and abnormal grain growth. Without loss of generality, the proposed phase-field model can be extended accordingly to include such anisotropy aspects.

### v3

@@ 3.3 Comparison to other grain growth phase-field models for oxide ceramics
$\eta_{i}$ phase-field order parameter of grain $i$, i=1,2,3,,ni=1,2,3,...,n -
$C_{\text{V}_{\text{O}}^{..}}$
Normalized local concentration of oxygen vacancy
($C_{\text{V}_{\text{O}}^{..}}=a^{3}c_{\text{V}_{\text{O}}^{..}}$, with $a$ being the lattice constant)
$C_{\text{dop}}$ Normalized local concentration of dopant, $C_{\text{dop}}=a^{3}c_{\text{dop}}$ -
$C_{\text{V}_{\text{O}}^{..},\text{b}}$ Normalized concentration of oxygen vacancy in bulk phase, $C_{\text{V}_{\text{O}}^{..},\text{b}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{b}}$ -
$C_{\text{V}_{\text{O}}^{..},\text{c}}$ Normalized concentration of oxygen vacancy in core region, $C_{\text{V}_{\text{O}}^{..},\text{c}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{c}}$ -
$C_{\text{dop},\text{b}}$ Normalized concentration of dopant in bulk phase, $C_{\text{dop},\text{b}}=a^{3}c_{\text{dop},\text{b}}$ -
$C_{\text{dop},\text{c}}$ Normalized concentration of dopant in core region, $C_{\text{dop},\text{c}}=a^{3}c_{\text{dop},\text{c}}$ -
@@
$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$
Formation energy of oxygen vacancy in the bulk phase.
$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}N_{\text{A}}$, with $N_{\text{A}}$ being Avogadro constant
$\frac{\text{J}}{\text{mol}}$
$g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$
Formation energy of oxygen vacancy in the core region.
$g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}N_{\text{A}}$
$\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{b}}$
Formation energy of dopant in the bulk phase.
$g^{0}_{\text{dop},\text{b}}=\mu_{\text{dop},\text{b}}^{0}N_{\text{A}}$
$\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{c}}$
Formation energy of dopant in the core region.
$g^{0}_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}N_{\text{A}}$
$\frac{\text{J}}{\text{mol}}$
@@
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$
Dimensionless available sites of oxygen vacancy in core region.
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}=N_{\text{V}_{\text{O}}^{..},\text{c}}a^{3}$, with $a$ being the lattice constant
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$
Dimensionless available sites of oxygen vacancy in bulk phase.
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}=N_{\text{V}_{\text{O}}^{..},\text{b}}a^{3}$
$\tilde{N}_{\text{dop},\text{c}}$
Dimensionless available sites of dopant in core region.
$\tilde{N}_{\text{dop},\text{c}}=N_{\text{dop},\text{c}}a^{3}$
$\tilde{N}_{\text{dop},\text{b}}$
Dimensionless available sites of dopant in bulk phase.
$\tilde{N}_{\text{dop},\text{b}}=N_{\text{dop},\text{b}}a^{3}$
@@ 4 Finite element implementation of the phase-field model
In this section, we present the detailed finite element implementation for the defect-chemistry informed phase-field model. According to the model presented in Section 3 and Table 2, we treat the OPs $\eta_{i}$ and the field variables $C_{\text{V}_{\text{O}}^{..}}$, $C_{\text{dop}}$, $C_{\text{V}_{\text{O}}^{..},\text{b}}$, $C_{\text{V}_{\text{O}}^{..},\text{c}}$, $C_{\text{dop},\text{b}}$, $C_{\text{dop},\text{c}}$, $\phi$ as degrees-of-freedom (DOFs). Moreover, to solve Eqs. (36) and (37), we use the mixed finite element formulation and treat the electro-chemical potentials $\mu_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}$ as additional DOFs bai2019two .
The strong forms of the corresponding governing equations have already been given in Eqs. (25), (26) and (34) to (40). According to the FEM hughes2012finite , we formulate the residuals from the weak forms of these governing equations by introducing corresponding test functions $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}$, $\psi_{C_{\text{dop},\text{c}}}$, $\psi_{\eta_{i}}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}$, $\psi_{C_{\text{dop}}}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}$, $\psi_{\mu_{\text{dop}}}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}$, $\psi_{C_{\text{dop},\text{b}}}$ and $\psi_{\phi}$. Firstly, the residuals for the weak forms of Eqs. (25), (26) and (34) are
@@
Following the Galerkin approach brenner2008mathematical , the test functions are discretized as $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}=N_{C_{\text{dop},\text{c}}}^{I}\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}=N_{\eta_{i}}^{I}\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}=N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}=N_{C_{\text{dop}}}\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}=N_{\mu_{\text{V}_{\text{O}}^{..}}}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}=N_{\mu_{\text{dop}}}\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}=N_{C_{\text{dop},\text{b}}}\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}=N_{\phi}\psi_{\phi}^{I}$, where $I$ denotes the node index and Einstein summation convention is used. $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}^{I}$ are the nodal weights, while $N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $N_{C_{\text{dop},\text{c}}}^{I}$, $N_{\eta_{i}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{C_{\text{dop}}}^{I}$, $N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{\mu_{\text{dop}}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $N_{C_{\text{dop},\text{b}}}^{I}$ and $N_{\phi}^{I}$ are the shape functions for the corresponding variables. Then we can write down the discretized residuals as
@@
$\begin{aligned} R_{\eta_{i}}^{I}&=\int_{V}\frac{N_{\eta_{i}}^{I}}{\Delta t}\text{d}V+L\int_{V}N_{\eta_{i}}^{I}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}N_{\eta_{i},j}^{I}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}N_{\eta_{i}}^{I}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}N_{\eta_{i}}^{I}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V\end{aligned}$ (57)
@@
The linearization of the residuals can be written in the element-level linear formulation, in which the terms in the tangent matrix are calculated by $K_{\xi,\zeta}^{IJ}=\partial R_{\xi}^{I}/\partial\zeta^{J}$, where $\xi,\zeta$} indicates individual DOF. The non-zero terms in the tangent matrix are presented in supplementary. The numerical implementation of the proposed phase-field model is carried out in Multiphysics Object-Oriented Simulation Environment (MOOSE) framework schwen2023phasefield , giudicelli2024moose . Spatial discretization was performed with linear elements for one-dimensional (1D) simulations (Section 5.1 and 5.2) and quadrilatera elements for two-dimensional (2D) simulations (Section 5.3), with linear Lagrange shape functions in each case aagesen2024electrochemical . The same basis functions were used for both trial (shape) and test functions, which is standard in the Galerkin finite element method brenner2008mathematical . Different mesh resolutions were adopted based on the specific requirements of each case. To ensure numerical accuracy, the relative convergence tolerance is set to $1\times 10^{-8}$. All phase-field simulations were conducted on CPU nodes equipped with Intel Xeon Platinum 8470Q processors and were accelerated using the Message Passing Interface (MPI) on the Lichtenberg high-performance computing cluster at Technische Universität Darmstadt.
@@ 5 Phase-field simulation results of Fe-doped SrTiO3
We present first in Section 5.1 one-dimensional (1D) phase-field simulation results to reproduce the formation of a SCL at equilibrium state under the influence of different defect chemistry parameters. Subsequently, we investigate the influence of solute drag effects induced by the segregation of point defects in the core on the formation of an asymmetric SCL under quasi-equilibrium state in Section 5.2. Finally, we conduct two-dimensional (2D) phase-field simulations to study the microstructure evolution process in STO polycrystalline material during sintering and analyze the resulting grain size distribution in Section 5.3.
Note that in the paper we ignore the GB orientation dependency or other anisotropy of the segregation energies $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$, for we intend to reveal the most fundamental mechanisms of SCL formation and the grain growth patterns. Without loss of generality, the proposed phase-field model can be extended accordingly to include such anisotropy aspects.